Use the Delta H degree_f and Delta H degree_rxn in
Answer #1

Hf = formation

Hrxn = HfSO3 - (HfSO2 + HfO2)

HfO2 = 0, by definition

then

Hrxn = -198

-198 = HfSO3 - (HfSO2 )

-198 = HfSO3 - -297

HfSO3 = -198-297 = -495 kJ/mol

which is answer c